OpenStudy (anonymous):

What is 0.004 written as a fraction of a percent? Review Comment For Answer

OpenStudy (anonymous):

A) 4/100 % B)1/4% C)1/3% D)2/5%

OpenStudy (anonymous):

0.004 = 4/1000 = 1/250 = 0.4% 0.4% is equivalent to 2/5% as a fraction. Therefore, the answer is d) 2/5%

OpenStudy (kymber):